# MEBench: Benchmarking Large Language Models for Cross-Document Multi-Entity Question Answering

### v1

@@ Abstract.
Multi-entity question answering (MEQA) demands the integration of scattered information across documents to resolve complex queries involving entities, relationships, and contextual dependencies. While large language models (LLMs) and retrieval-augmented generation (RAG) systems show promise, their performance on cross-document MEQA remains underexplored due to the absence of tailored benchmarks. To address this gap, we introduce MEBench, a novel multi-document, multi-entity benchmark designed to systematically evaluate LLMs’ capacity to retrieve, consolidate, and reason over scattered and dense information. Our benchmark comprises 4,780 questions which are systematically categorized into three primary categories: Comparative Reasoning, Statistical Reasoning and Relational Reasoning, further divided into eight distinct types, ensuring broad coverage of real-world multi-entity reasoning scenarios. Our experiments on state-of-the-art LLMs reveal critical limitations: even advanced models achieve only 59% accuracy on MEBench. Our benchmark emphasizes the importance of completeness and factual precision of information extraction in MEQA tasks, using Entity-Attributed F1 (EA-F1) metric for granular evaluation of entity-level correctness and attribution validity. MEBench not only highlights systemic weaknesses in current LLM frameworks but also provides a foundation for advancing robust, entity-aware QA architectures.
@@
Multi-entity question answering, Large language models (LLMs), Cross-document reasoning
doi: conference: ACM SIGKDD Conference on Knowledge Discovery and Data Mining Datasets and Benchmarks Track; August 03–07, 2025; Toronto, Canadaisbn: ccs: Information systems Question answering
1. Introduction
@@
The emergence of large language models (LLMs) has significantly advanced natural language processing capabilities, demonstrating exceptional performance in diverse tasks spanning text generation to complex logical reasoning (Achiam et al., 2023) (Meta Llama3, 2024). Nevertheless, their effectiveness in addressing cross-document multi-entity question answering (MEQA), a task requiring the integration of fragmented, entity-specific information across heterogeneous document sources, remains insufficiently investigated. Current implementations of retrieval-augmented generation (RAG) architectures (Wu et al., 2025) (Fan et al., 2024) (Tang et al., 2024) and long-context LLM frameworks exhibit notable limitations in processing entity-dense analytical reasoning, particularly when contextual dependencies are distributed across multiple documents, and we analytically argue that context window limitations, over-reliance on parametric knowledge, and poor cross-document attention as the key bottlenecks. Furthermore, the field lacks comprehensive benchmarking frameworks specifically designed to evaluate cross-document entity-intensive scenarios (Song et al., 2024). As shown in Figure 1, existing evaluation metrics frequently inadequately represent the complexities inherent in real-world MEQA applications, where queries such as “What is the distribution of Turing Award winners by fields of study?” necessitate not only high-precision information retrieval but also structured synthesis of overlapping, contradictory, or complementary data points extracted from disparate documents.
To address this methodological gap, we present MEbench, a novel benchmarking framework specifically designed to assess the performance of large language models and RAG systems in cross-document multi-entity question answering scenarios. The benchmark simulates real-world information integration challenges where correct answers require synthesizing entity-centric evidence distributed across multiple documents, with a single instance of document omission or entity misinterpretation can propagate errors through the reasoning chain.
@@
As shown in Table 2, MEBench features a mean entity density of 409 entities per query, with systematically varied entity cardinality across three operational tiers: low (0-10 entities), medium (10-100 entities), and high complexity (¿100 entities). This stratified design enables granular performance evaluation across different entity scales and task difficulty levels. The framework comprises 4,780 validated question-answer pairs systematically categorized into three primary categories and eight distinct types, MEBench spans diverse real-world scenarios, from academic field distributions to geopolitical event analysis.
Our experiments with state-of-the-art models, including GPT-4 and Llama-3, reveal significant shortcomings: even the most advanced LLMs achieve only 59% accuracy on MEBench. This underscores systemic weaknesses in current frameworks, for example, models frequently fail to locate all entity attributes or infer implicit relationships, highlighting the need for architectures that prioritize entity-aware retrieval and contextual consolidation.
Our main contributions are summarized as follows:
@@
  • Development of MEBench: A novel multi-document, multi-entity benchmark designed to evaluate LLMs and RAG systems in cross-document aggregation and reasoning. It includes 4,780 validated question-answer pairs spanning three categories and eight types, simulating real-world scenarios that demand integration of fragmented, entity-centric information.
@@
  • Entity-centric Task Categories and Evaluation: Utilization of Entity-Attributed F1 (EA-F1), a granular metric for assessing entity-level correctness and attribution validity, alongside a stratified entity density design (low: 0–10, medium: 11–100, high: ¿100 entities per query). Our framework emphasizes completeness and factual precision in information extraction, addressing gaps in existing metrics for entity-dense MEQA tasks.
@@
  • Scalable Benchmark Construction. A scalable, automated pipeline featuring: Knowledge graph extraction from structured Wikipedia for cross-document relationship discovery; Relational table generation to preserve entity-property relationships; Template-based QA generation ensuring reproducibility and reducing cost and labor.
@@
Figure 1. Existing benchmarks vs. MEBen. Unlike existing benchmarks which feature centralized evidence distributions and sparse entity mentions, MEBen presents entity-dense scene where critical evidences are dispersed across multiple documents, necessitating that when seeking an answer, no document or entity can be ignored.
2. Related Work
@@
2.1. Single-Document QA and LLM Progress.
@@
Many QA benchmarks, such as SQuAD (Rajpurkar et al., 2016), Natural Questions  (Kwiatkowski et al., 2019), L-eval (An et al., 2024) and needle-in-a-haystack (Kamradt, 2023), focus on extracting answers from individual document. Modern LLMs like GPT-4 (Achiam et al., 2023), Llama-3 (Meta Llama3, 2024), and PaLM (Chowdhery et al., 2023) have achieved near-human performance on these tasks, leveraging their ability to parse and reason within localized contexts. However, these benchmarks do not address the complexities of integrating information across multiple documents, a critical limitation for real-world applications.
2.2. Cross-Document Aggregation Challenges.
@@
Efforts to extend QA to multi-document settings include datasets like HotpotQA (Yang et al., 2018), MuSiQue (Trivedi et al., 2021), LooGLE (Li et al., 2024), LM-Infinit (Han et al., 2024), $\infty$ Bench (Zhang et al., 2024), CLongEval (Qiu et al., 2024), BAMBOO (Dong et al., 2024) and Loong (Wang et al., 2024), which emphasize multi-hop reasoning and cross-source synthesis. While these benchmarks highlight the need for systems to connect disparate information, they often prioritize breadth over depth in entity-centric reasoning. For instance, questions in these datasets rarely demand the consolidation of attributes for dozens of or more entities (e.g., aggregating ACM Fellows’ expertise across fields), a gap that limits their utility in evaluating entity-dense scenarios. Recent RAG frameworks (Fan et al., 2024) aim to enhance retrieval-augmented QA but struggle with ensuring completeness and attribution validity when handling multi-entity queries.
2.3. Entity-Centric Evaluation Metrics.
@@
Existing evaluation metrics for QA, such as F1 score and exact match (EM), focus on answer surface-form correctness but overlook granular entity-level attribution (Rostampour et al., 2010). Metrics in FEVER (Thorne et al., 2018) and Attributed QA (Bohnet et al., 2023) emphasize source verification, yet they lack the specificity to assess multi-entity integration. For example, they do not systematically measure whether all relevant entities are retrieved, their attributes are correctly extracted, or their sources are accurately used, which is a shortcoming that becomes critical in MEQA tasks.
2.4. The Gap in Multi-Entity QA Benchmarks.
@@
Prior work has yet to establish a benchmark that systematically evaluates LLMs and RAG systems on entity-dense, cross-document reasoning. Current datasets either lack the scale and diversity of real-world multi-entity questions or fail to provide fine-grained metrics for assessing entity-level completeness and attribution (Wang et al., 2024) (Bai et al., 2025). MEBench addresses these limitations by introducing a comprehensive evaluation framework that challenges models to retrieve, consolidate, and reason over scattered entity-centric data across heterogeneous sources. By incorporating the Entity-Attributed F1 (EA-F1) metric, our benchmark advances the field toward more precise, entity-aware QA systems.
3. MEBench
Table 1. Examples of multi-entities queries.
@@ 3. MEBench
Table 2. Statistics of MEBench benchmark.
@@
#-high (¿100) 1946 785 2731
Figure 2. The systematic pipeline of Benchmark Construction. It comprising three phases: documents collection, information extraction and question-answer generation. In the documents collection phase, concept topics relevant to multi-entity scenarios are selected, followed by GPT-4 processing descriptions to extract entities and properties mapped to Wikipedia IDs for integration with structured Wiki data. Structured information from Wikipedia documents is processed using small language models (SLMs) due to the structured nature of the documents, culminating in table creation with entity attributes as columns. For QA generation, questions are generated following a ”template-driven, entity-attribute coupling” paradigm using GPT-4 with predefined templates, and undergo syntactic, semantic, and ambiguity checks, while answers are programmatically derived via SQL queries against the table and standardized into canonical forms. The final dataset ensures traceability (SQL-linked answers), scalability (template-driven approach), and rigor (execution-based answering reduces hallucination risks).
3.1. Task overview
MEBench is a structured evaluation framework designed to systematically assess the capabilities of LLMs in performing cross-document multi-entity question answering. This framework targets three core reasoning modalities: comparative analysis, statistical inference, and relational reasoning, and each decomposed into specialized subtasks that rigorously test distinct facets of LLM performance. (Examples are provided in Table  1), ensuring broad coverage of real-world multi-entity reasoning scenarios. MEBench evaluates LLMs on cross-document MEQA through three primary task categories, each addressing distinct reasoning challenges:
3.1.1. Comparative Reasoning
@@
Intercomparison (Cross-Document Entity Contrast)
: Intercomparison task requires models to perform multi-faceted comparisons of attributes or relationships among entities distributed across documents. This involves not only identifying relevant information from disparate sources but also aligning it in a coherent manner to facilitate direct comparison.
Superlative Identification
: This subtask tests the model’s capacity to identify maxima/minima or rank entities based on quantitative or qualitative criteria.This task tests the model’s proficiency in quantitative analysis and its ability to handle superlative queries that require precision and attention to detail.
3.1.2. Statistical Reasoning
@@
Statistical tasks (Zhu et al., 2024) assess LLM’s proficiency in quantitative synthesis, including aggregation, distributional analysis, correlation analysis, and variance analysis across multi-document.
Aggregation
: For aggregation task, models must calculate cumulative values from data spread across different documents, demonstrating its ability to perform basic arithmetic operations and handle aggregation queries effectively.
Distribution Compliance
: Here, the focus is on understanding and analyzing statistical distributions. This task requires the model to parse data, categorize them appropriately, and compute proportions, thereby showcasing its ability to comprehend and manipulate statistical data.
Correlation Analysis
: Tests the ability of LLM to infer relationships between variables. This task tests the LLM’s capability to understand complex relationships and draw logical inferences based on evidence from multiple sources.
Variance Analysis (Outlier Detection)
: Assesses sensitivity to data variability and anomalies. This requires the model to perform statistical analysis to understand variability and its implications.
3.1.3. Relational Reasoning
@@
Relational tasks probe an LLM’s capacity to model explicit interactions and counterfactual dependencies among entities.
Descriptive Relationship
:Examines the model’s ability to reconstruct explicit entity interactions. Descriptive relationship tests the model’s ability to construct coherent narratives based on factual data.
Hypothetical Scenarios
: This task evaluates LLM’s ability to engage in counterfactual reasoning, a critical component of causal inference and speculative analysis. Model must simulate alternate realities by removing or altering specific entities or events and hypothesizing downstream impacts, demonstrating its ability to engage in abstract and speculative thinking.
3.2. Benchmark Construction
@@
MEBench was constructed through a systematic pipeline:
3.2.1. Data Collection
Concept Topic Identification
@@ 3.2.1. Data Collection
. In the initial phase of data collection for MEbench, a meticulous process is employed to determine the concept topics that are applicable to multi-entity scenarios. These topics are carefully selected based on their significance, prevalence, and the potential for generating complex multi-entity questions, and examples can be seen in Appedix Table 5.
Entity and Property Identification
@@
. Once the concept topics are determined, We input descriptions related to the concept topics into the LLM (GPT-4), which then processes the text to identify concept entities and property, as illustrated in Figure 2-a1. After the LLM identifies the entities and Property via iterative semantic refinement, we map them to entity IDs and Property IDs in the Wiki graph. This mapping is crucial as it allows for seamless integration with the vast amount of structured data available in Wikipedia. The detailed method is in Appendix A.1. Using the Entity ID and property ID, we synthesise SPARQL. We then utilize the API provided by Wikipedia to retrieve the wiki web pages of all entities related to the topic. For example, if our concept topic is ”ACM Fellows”, we would obtain the Wikipedia pages of all ACM Fellows, which contain their detailed information. We use GPT-4 to generate a set of interesting entity attributes. These attributes are carefully chosen based on the general interest and relevance in the domain. For ACM Fellows as example, nationality, research field, institution, and academic contribution are attributes that people commonly pay attention to.
Structured Information Processing
@@
. Once the document set is collected, we proceed to the structured information processing stage. The documents we have gathered from Wikipedia have well-defined and accurate structural relations. Due to the structured nature of the documents, we do not need to rely on the long context ability of large language models. Instead, we can use small language models (SLMs) for information extraction. They are well-suited for tasks where the information is already structured and the focus is on extracting specific details (Fan et al., 2025).
Table Generation
@@
. The final step in the data collection process is to generate a table, as shown in Figure 2-b1. We use the extracted information and the entity attributes as the column headers of the table. Each row in the table represents an individual entity. For example, in the case of ACM Fellows, each row would correspond to an individual ACM Fellow.
3.2.2. Question and answer Generation

### v2

@@ Abstract
Cross-Document Multi-entity Question Answering (MEQA) demands the integration of scattered information across documents to resolve complex queries involving entities, relationships, and contextual dependencies. Although Large Language Models (LLMs) and Retrieval-augmented Generation (RAG) systems show promise, their performance on cross-document MEQA remains underexplored due to the absence of tailored benchmarks. To address this gap, we introduce MEBench, a scalable multi-document, multi-entity benchmark designed to systematically evaluate LLMs’ capacity to retrieve, consolidate, and reason over scattered and dense information. Our benchmark comprises 4,780 questions which are systematically categorized into three primary categories: Comparative Reasoning, Statistical Reasoning and Relational Reasoning, further divided into eight distinct types, ensuring broad coverage of real-world multi-entity reasoning scenarios. Our experiments on state-of-the-art LLMs reveal critical limitations: even advanced models achieve only 59% accuracy on MEBench. Our benchmark emphasizes the importance of completeness and factual precision of information extraction in MEQA tasks, using Entity-Attributed F1 (EA-F1) metric for granular evaluation of entity-level correctness and attribution validity. MEBench not only highlights systemic weaknesses in current LLM frameworks but also provides a foundation for advancing robust, entity-aware QA architectures. [^1] The source code and data have been made available at https://github.com/tl2309/MEBench
@@
MEBench: Benchmarking Large Language Models for Cross-Document Multi-Entity Question Answering
Teng Lin1, Yuyu Luo1,2, Nan Tang1,2 1The Hong Kong University of Science and Technology (Guangzhou) 2The Hong Kong University of Science and Technology tlin280@connect.hkust-gz.edu.cn {yuyuluo, nantang}@hkust-gz.edu.cn
1 Introduction
@@
The emergence of large language models (LLMs) has significantly advanced natural language processing capabilities, demonstrating exceptional performance in diverse tasks spanning text generation to data science and databases Achiam et al. (2023); Lin (2025); Liu et al. (2025a); Li et al. (2025a); Zhang et al. (2025); Li et al. (2024a); Chen et al. (2025); Li et al. (2025b); Fan et al. (2024a). Nevertheless, long-context LLMs exhibit notable limitations in processing entity-dense analytical reasoning, particularly when contextual dependencies are distributed across multiple documents Wu et al. (2025c, a); Shi et al. (2025), and we analytically argue that context window limitations, over-reliance on parametric knowledge, and poor cross-document attention as the key bottlenecks Tang et al. (2024b); Zhu et al. (2024); Xiang et al. (2025); Lin et al. (2025b). On the other hand, current implementations of retrieval-augmented generation (RAG) architectures Yang et al. (2024); Wu et al. (2025b); Fan et al. (2024b); Tang et al. (2024a); Liu et al. (2025b); Zhang et al. (2024b); Lin et al. (2025a); Hong et al. (2025) frameworks’ effectiveness in addressing cross-document Multi-entity Question Answering (MEQA) remains insufficiently investigated. Furthermore, the field lacks comprehensive benchmarking frameworks specifically designed to evaluate the performance of LLMs and RAG systems for cross-document entity-intensive tasks. As shown in Figure 1, existing evaluation metrics frequently inadequately represent the complexities inherent in real-world MEQA applications (Song et al., 2024), where queries such as “What is the number distribution of all Turing Award winners by fields of study by 2023?” necessitate not only high-precision information retrieval but also reasoning over fragmented, entity-specific information across heterogeneous document sources.
@@
To address this methodological gap, we present MEbench, a novel benchmarking framework specifically designed to assess the performance of large language models and RAG systems in cross-document multi-entity question answering scenarios. The benchmark simulates real-world information integration challenges where correct answers require synthesizing entity-centric evidence distributed across multiple documents, with a single instance of document omission or entity misinterpretation can propagate errors through the reasoning chain. As shown in Table 2, MEBench features a mean entity density of 409 entities per query, with systematically varied entity cardinality across three operational tiers: low (0-10 entities), medium (10-100 entities), and high complexity (>100 entities). This stratified design enables granular performance evaluation across different entity scales and task difficulty levels. The framework comprises 4,780 validated question-answer pairs systematically categorized into three primary categories and eight distinct types, MEBench spans diverse real-world scenarios, from academic field distributions to geopolitical event analysis. Our experiments with state-of-the-art models, including GPT-4 and Llama-3, reveal significant shortcomings: even the most advanced LLMs achieve only 59% accuracy on MEBench. This underscores systemic weaknesses in current frameworks, for example, models frequently fail to locate all entity and their attributes or infer implicit relationships, highlighting the need for architectures that prioritize entity-aware retrieval and contextual consolidation.
Our contributions are summarized as follows:
@@
Development of MEBench. A scalable benchmark designed to evaluate LLMs and RAG systems in cross-document aggregation and reasoning. It includes 4,780 validated question-answer pairs spanning three categories and eight types, simulating real-world scenarios that demand integration of scattered, entity-specific information.
@@
Entity-centric Task Categories and Evaluation. Utilization of Entity-Attributed F1 (EA-F1), a granular metric for assessing entity-level correctness and attribution validity, alongside a stratified entity density design (low: 0–10, medium: 11–100, high: >100 entities per query). Our framework emphasizes completeness and factual precision in information extraction, addressing gaps in existing metrics for entity-dense MEQA tasks.
@@
Scalable Benchmark Construction. A scalable, automated pipeline: Knowledge graph extraction from structured Wikipedia for cross-document relationship discovery; Relational table generation to preserve entity-property relationships; Template-based QA generation ensuring reproducibility and reducing cost and labor.
@@
Figure 1: Existing benchmarks vs. MEBench. Unlike existing benchmarks which feature centralized evidence distributions and sparse entity mentions, MEBench presents entity-dense scene where critical evidences are dispersed across multiple documents, necessitating that when seeking an answer, no document or entity can be ignored.
2 Related Work
@@
2.1 Single-Document QA and LLM Progress
@@
Many QA benchmarks, such as SQuAD Rajpurkar et al. (2016), Natural Questions  Kwiatkowski et al. (2019), L-eval An et al. (2024) and needle-in-a-haystack Kamradt (2023), focus on extracting answers from individual document. Modern LLMs like GPT-4 Achiam et al. (2023), Llama-3 Meta Llama3 (2024), and PaLM Chowdhery et al. (2023) have achieved near-human performance on these tasks, leveraging their ability to parse and reason within localized contexts. However, these benchmarks do not address the complexities of integrating information across multiple documents, a critical limitation for real-world applications Luo et al. (2018a, b, 2021, 2022); Qin et al. (2020); Zhu et al. (2025); Liu et al. (2025c, a).
2.2 Cross-Document Aggregation Challenges
@@
Efforts to extend QA to multi-document settings include datasets like HotpotQA Yang et al. (2018), MuSiQue Trivedi et al. (2021), LooGLE Li et al. (2024b), LM-Infinit Han et al. (2024), $\infty$ Bench Zhang et al. (2024a), CLongEval Qiu et al. (2024), BAMBOO Dong et al. (2024), Loong Wang et al. (2024) and Symphony Chen et al. (2023), which emphasize multi-hop reasoning and cross-source synthesis. While these benchmarks highlight the need for systems to connect disparate information, they often prioritize breadth over depth in entity-centric reasoning. For instance, questions in these datasets rarely demand the consolidation of attributes for dozens or more entities (e.g., aggregating ACM Fellows’ expertise across fields), a gap that limits their utility in evaluating entity-dense scenarios. Recent RAG frameworks Fan et al. (2024b); DBLP:journals/corr/abs-2504-10036 aim to enhance retrieval-augmented QA but struggle with ensuring completeness and attribution validity when handling multi-entity queries.
2.3 Entity-Centric Evaluation Metrics.
@@
Existing evaluation metrics for QA, such as F1 score and exact match (EM), focus on answer surface-form correctness but overlook granular entity-level attribution Rostampour et al. (2010). Metrics in FEVER Thorne et al. (2018), Attributed QA Bohnet et al. (2023) and emphasize source verification, yet they lack the specificity to assess multi-entity integration. For example, they do not systematically measure whether all relevant entities are retrieved, their attributes are correctly extracted, or their sources are accurately used, which is a shortcoming that becomes critical in MEQA tasks.
2.4 The Gap in Multi-Entity QA Benchmarks.
@@
Prior work has yet to establish a benchmark that systematically evaluates LLMs and RAG systems on entity-dense, cross-document reasoning. Current datasets either lack the scale and diversity of real-world multi-entity questions or fail to provide fine-grained metrics for assessing entity-level completeness and attribution Song et al. (2024); Wang et al. (2024); Bai et al. (2025). MEBench addresses these limitations by introducing a comprehensive evaluation framework that challenges models to retrieve, consolidate, and reason over scattered entity-centric data across heterogeneous sources. By incorporating the Entity-Attributed F1 (EA-F1) metric, our benchmark advances the field toward more precise, entity-aware QA systems.
3 MEBench
Table 1: Examples of multi-entities queries.
@@ 3 MEBench
Table 2: Statistics of MEBench benchmark.
@@
#-high (>100) 1946 785 2731
3.1 Task overview
MEBench is a structured evaluation framework designed to systematically assess the capabilities of LLMs and RAG systems in performing cross-document multi-entity question answering. This framework targets three core reasoning modalities: comparative analysis, statistical inference, and relational reasoning, and each decomposed into specialized subtasks that rigorously test distinct facets of LLM performance, ensuring broad coverage of real-world multi-entity reasoning scenarios. Examples of tasks are provided in Table  1. Each of three primary task categories addresses distinct reasoning challenges:
Comparative Reasoning
@@ 3.1 Task overview
Statistical Reasoning
@@
Statistical tasks assess LLM’s proficiency in quantitative synthesis, including aggregation, distributional analysis, correlation analysis, and variance analysis across multi-document.
Relational Reasoning
@@
Relational tasks probe model’s capacity to model explicit interactions and counterfactual dependencies among entities.
3.2 Benchmark Construction
@@
MEBench was constructed through a systematic pipeline, comprising the following steps.
Figure 2: The systematic pipeline of benchmark construction. It comprising three phases: documents collection, information extraction and question-answer generation. In the documents collection phase, concept topics relevant to multi-entity scenarios are selected, followed by GPT-4 processing descriptions to extract entities and properties mapped to Wikipedia IDs for integration with structured Wiki data. Structured information from Wikipedia documents is processed using small language models (SLMs) due to the structured nature of the documents, culminating in table creation with entity attributes as columns. For QA generation, questions are generated following a "template-driven, entity-attribute coupling" paradigm using GPT-4 with predefined templates, and undergo syntactic, semantic, and ambiguity checks, while answers are programmatically derived via SQL queries against the table and standardized into canonical forms. The final dataset ensures traceability (SQL-derived answers), scalability (template-driven approach), and rigor (execution-based answering reduces hallucination risks).
3.2.1 Data Collection
Concept Topic Identification.
@@ 3.2.1 Data Collection
In the initial phase of data collection for MEbench, a meticulous process is employed to determine the concept topics that are applicable to multi-entity scenarios. These topics are carefully selected based on their significance, prevalence, and the potential for generating complex multi-entity questions, and examples can be seen in Appendix Table 5.
Entity and Property Identification.
@@
Once the concept topics are determined, we input descriptions related to the concept topics into a LLM (we use GPT-4), which then processes the text to identify concept entity and property, as illustrated in Figure 2-a1. After the LLM identifies the entity and property via iterative semantic refinement, we map them to entity IDs and property IDs in the Wiki graph. This mapping is crucial as it allows for seamless integration with the vast amount of structured data available in Wikipedia. The detailed method is in Appendix A.1. Using the entity ID and property ID, we synthesise SPARQL. We then utilize the API provided by Wikipedia to retrieve the wiki web pages of all entities related to the topic. For example, if our concept topic is "ACM Fellows", we would obtain the Wikipedia pages of all ACM Fellows, which contain their detailed information. We use GPT-4 to generate a set of interesting entity attributes. These attributes are carefully chosen based on general interest and relevance in the domain. For ACM Fellows, as an example, nationality, research field, institution, and academic contribution maybe the attributes that people commonly pay attention to.
Structured Information Processing.
@@
Once the document set is collected, we proceed to the structured information processing stage. The documents we have gathered from Wikipedia have well-defined and accurate structural relations. Due to the structured nature of the documents, we do not need to rely on the long context ability of large language models. Instead, we can use small language models (SLMs) for information extraction. They are well-suited for tasks where the information is already structured and the focus is on extracting specific details (Fan et al., 2025).
Table Generation.
@@
The final step in the data collection process is to generate a table, as shown in Figure 2-b1. We use the the entity attributes as the column headers of the table. Each row in the table represents an individual entity. For example, in the case of ACM Fellows, each row would correspond to an individual ACM Fellow.
3.2.2 Question and answer Generation